OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, 3) and perpendicular to y = -1/3 x + 7.

OpenStudy (anonymous):

@Elsa213 can you help please

Elsa213 (elsa213):

Two lines are perpendicular if their slopes are opposite recipricals. That means that the equation to compare the slopes is: m1 = -1/m2, where m1 and m2 are the slopes of the two lines. The equation for the line that is perpendicular to the line we want is y =-1/4x -5, which means its slope is -1/4. m1 = -1/(-1/4) Equation for perpendicular slope with -1/4 substituted for m2 m1 = 4 Simplified Now that we have the slope of the line that we want (4) and a point (-1,7), we can use point-slope formula to find the equation of the line that we want. y - y1 = m(x - x1) Point-slope formula y - 7 = 4(x - (-1)) Substitution y - 7 = 4(x + 1) Subtracting a negative means adding a positive y - 7 = 4x + 4 Distributive property y - 7 + 7 = 4x + 4 + 7 Add 7 to each side y = 4x + 11 Simplify

OpenStudy (anonymous):

I really dont get it can you explain